# Preprint submitted to Mathematical Geosciences

### v1

@@ Abstract
Geological parameterization procedures entail the mapping of a high-dimensional geomodel to a low-dimensional latent variable. These parameterizations can be very useful for history matching because the number of variables to be calibrated is greatly reduced, and the mapping can be constructed such that geological realism is automatically preserved. In this work, a parameterization method based on generative latent diffusion models (LDMs) is developed for 3D channel-levee-mud systems. Geomodels with variable scenario parameters, specifically mud fraction, channel orientation, and channel width, are considered. A perceptual loss term is included during training to improve geological realism. For any set of scenario parameters, an (essentially) infinite number of realizations can be generated, so our LDM parameterizes over a very wide model space. New realizations constructed using the LDM procedure are shown to closely resemble reference geomodels, both visually and in terms of one- and two-point spatial statistics. Flow response distributions, for a specified set of injection and production wells, are also shown to be in close agreement between the two sets of models. The parameterization method is applied for ensemble-based history matching, with model updates performed in the LDM latent space, for cases involving geological scenario uncertainty. For three synthetic true models corresponding to different geological scenarios, we observe clear uncertainty reduction in both production forecasts and geological scenario parameters. The overall method is additionally shown to provide posterior geomodels consistent with the synthetic true model in each case.
Keywords: Geological parameterization, diffusion models, latent diffusion, history matching, scenario uncertainty, multiple scenario
@@ 1 Introduction
In this work, we introduce a 3D deep learning-based latent diffusion model for the parameterization of facies-based geological systems. We consider relatively large channel-levee-mud geomodels, and introduce a parameterization that can treat multiple geological scenarios characterized by (variable) metaparameters. This represents a significant extension of our previous latent diffusion model where we considered 2D, single-scenario systems (Di Federico and Durlofsky, 2025). We demonstrate the usefulness of the parameterization method by performing history matching with an ensemble smoother with multiple data assimilation (ESMDA) procedure (Emerick and Reynolds, 2013) operating on latent space variables.
@@
In recent years, a number of deep learning (DL) geomodel parameterizations have been introduced. These have been shown to perform better than more mathematically tractable methods, such as those based on principal component analysis (e.g., Sarma et al. (2008)) or discrete cosine transform (Jafarpour and McLaughlin, 2009), which are not always able to reproduce complex geological patterns (Chan and Elsheikh, 2019). DL methods learn the mapping between geological model space and latent space from a training dataset, and are then able to generate new realizations that exhibit the same spatial features. Architectures for this application include CNN-PCA (Liu and Durlofsky, 2021), variational autoencoders (VAEs) (Canchumuni et al., 2019; Mo et al., 2020; Liu and Grana, 2020), and generative adversarial networks (GANs) (Laloy et al., 2018; Zhang et al., 2019; Canchumuni et al., 2021; Hu et al., 2023; Feng et al., 2024b; Merzoug and Pyrcz, 2025). Existing DL parameterizations have proved successful both for standalone geomodeling and history matching for various subsurface flow applications. These include, for example, oil and gas production (Tang et al., 2021), groundwater flow (Song et al., 2021c), and geological carbon storage (Bao et al., 2024). These methods do, however, have some limitations. Namely, VAEs may provide overly smoothed or discontinuous geological patterns (Canchumuni et al., 2019; Zhang et al., 2019; Abbate et al., 2024), while GANs often require computationally intensive (potentially unstable) adversarial training in conjunction with careful hyperparameter tuning. Improved treatments have been presented, including VAE-GAN hybrids (Zhao and Chen, 2023; Feng et al., 2024a), progressive training (Song et al., 2021c), StyleGAN (Ling and Jafarpour, 2024), and the use of the Wasserstein loss (Chan and Elsheikh, 2019). Although these developments are promising, the limitations of existing methods have motivated the application of the more recent diffusion models (DMs), which currently represent the state-of-the-art in generative modeling.
@@
DMs, first introduced by  Ho et al. (2020), are a class of generative models trained to learn a denoising process. During denoising, a neural network is used to predict the noise contained in a sample, which is then iteratively subtracted for a number of steps. In this way, the DM transforms an initial noise input into a clean, “denoised” image. To accomplish this, the DM learns the mapping between a Gaussian noise realization (latent space) and the corresponding geological model (model space). DMs have been applied for geomodeling purposes in a few studies, and they have been shown to provide high-quality unconditional and conditional realizations for complex geological patterns. These include fluvial channel (Mosser, 2023; Xu et al., 2024; Di Federico and Durlofsky, 2025), Gaussian (Alnasser and Maucec, 2024), and discrete fracture systems (Teng et al., 2024).
@@
Among DMs, latent diffusion models (LDMs, introduced by Rombach et al. (2022)), appear to be a particularly promising method for geomodel generation and history matching in subsurface flow applications. These methods achieve computational efficiency and scalability by performing the diffusion process in a lower-dimensional latent space rather than in the pixel or voxel space.  Lee et al. (2025) and  Ovanger et al. (2025) used LDMs to generate realistic conditional realizations of 2D multifacies marine depositional environments. These studies highlighted the superior performance of diffusion-based generation relative to GANs. In Teng et al. (2024), an LDM was used to parameterize 2D discrete fracture networks (DFNs). LDM generated models exhibited visual improvement over those generated using VAEs, and they preserved property distributions in agreement with reference models. LDMs have also been used for direct generation or post-processing of simulation forecasts. For example, Zhan et al. (2025) applied an LDM for 2D binary facies models with Gaussian permeability distributions within the two facies. In that study an LDM was trained to generate facies structures conditioned on observed pressure and concentration fields, and another LDM with the same architecture was trained to perform the opposite task. Kim et al. (2025) developed an LDM model to post-process upscaled simulation results to provide fine-scale pressure and saturation maps for a 2D waterflooding problem.
@@
To date, many subsurface modeling applications involving DL parameterizations – and in particular most DM/LDM implementations – have been for 2D or relatively small 3D domains, containing fewer than $\sim$100,000 cells. Given their powerful generative capabilities and efficient representation of high-dimensional spaces, the extension and application of these methods to larger 3D models is of great interest. In addition, with the exception of the work in Song et al. (2021b) for 2D geomodeling with GANs, only a single geological scenario has been considered in most studies involving flow and history matching. This also represents a key limitation because the geological scenario itself is typically uncertain and this should be considered during history matching.
@@
In this work, we address both of these important issues by (1) developing a multiscenario 3D-LDM capability, and (2) applying this treatment to history matching for cases where the geological scenario is uncertain. The 3D models considered in this work, of grid dimensions $128\times 128\times 32$ (total of 524,288 cells), are larger than many of the models treated in previous studies. In addition, the three-facies geomodels entail uncertain scenario parameters, namely mud fraction, channel width, and channel orientation.
@@
New samples generated by the 3D-LDM method are shown to exhibit high visual quality in terms of geological features, along with spatial statistics that are consistent with (reference) training samples from Petrel geomodeling software (Schlumberger, 2019). Significant computational savings are achieved in the geomodeling process. Flow statistics for prior models confirm that the parameterization provides median results, and the correct amount of variability, relative to reference results across geological scenarios. History matching is then performed for waterflooding problems using the 3D-LDM representation in combination with an ESMDA procedure (Emerick and Reynolds, 2013). Results for cases involving synthetic true models from different geological scenarios are presented.
@@
This paper proceeds as follows. In Section 2, we describe the channel-levee-mud geomodels considered in this work. In Section 3, after reviewing the basics of DMs/LDMs and their use for geomodeling, we describe the 3D-LDM parameterization method and the new treatments introduced in this work. Next, in Section 4, we display LDM geomodel realizations and compare spatial and flow statistics for an ensemble of LDM generated models to those for reference Petrel models. Latent-space history matching results, for different synthetic true models, are provided in Section 5. Finally, in Section 6, we summarize this work and suggest directions for future research. Some algorithmic details appear in the Appendix.
2 Multiple scenario 3D fluvial channel systems
@@
In this work, we define a geological scenario as a realization of a set of uncertain geological scenario parameters (or metaparameters) that characterize the depositional environment. This approach allows for variation within a general geological concept. For example, in the systems considered here, the metaparameters quantify the geometries of channel bodies in fluvial channel systems. It is important to emphasize that, with the scenario parameters considered here, we are not attempting to treat widely different depositional environments corresponding to distinct geological concepts. For example, these parameters are not meant to characterize settings ranging from fluvial channel systems to eolian sand-dune systems to fractured carbonate systems. Nonetheless, even within a particular geological concept, an infinite number of realizations can be constructed for any set of scenario parameters. In this sense the geomodel uncertainty is hierarchical, as we have uncertain higher-level scenario parameters and, within each scenario, uncertain geobody realizations.
@@
The fluvial channel systems considered here contain channel (high permeability), levee (intermediate permeability), and mud/shale (low permeability) facies. The levee facies surrounds the extensive channel bodies, with both deposited within a mud background. Mud, levee, and channel facies are identified in the geomodels by cell values of 0, 0.5, and 1, respectively. The geological scenario parameters include mud fraction, channel orientation, and channel width. The geomodel realizations used for training, and as the reference against which the LDM models will be compared, are constructed using the Petrel geomodeling software (Schlumberger, 2019). Object-based modeling (OBM) is applied.
@@
Mud fraction 0.72 0.87
Channel orientation (avg. angle) 30 60
Channel width (avg. # of blocks) 4 6
@@
Figure 1: (a) Conditioning locations for all geomodels and well labels for (later) flow simulations. (b–f) Representative realizations corresponding to different geological scenarios from the Petrel-generated training set (scenario parameters are given for each geomodel, channel width in units of grid cells).
@@ 3 Latent diffusion models for representing geological systems
In this section, we first present an overview of the general DM (Ho et al., 2020) and LDM (Rombach et al., 2022) procedures. Our 3D-LDM method for multiscenario geological parameterization is then described. From a geomodeling perspective, scenario parameters act as a form of conditioning to global features, as opposed to conditioning to particular values at specific spatial locations, i.e., hard data. Strategies to implement scenario conditioning in a deep learning parameterization context include the use of additional input channels or separate parameter-encoding pipelines. These approaches enable direct control on the generation process, and have proved effective for geomodeling purposes (Song et al., 2021c). In this study, however, we adopt a simpler approach, where the method is trained to parameterize over a range of scenarios without additional constraints or input features.
@@ 3.1 Diffusion model (DM)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\mathbf{I}_{N_{c}}$ is the identity matrix. The scheduler defines the values of the perturbation through the $\alpha_{t}$ parameter at each step. This is typically a linearly decreasing function from $\alpha_{1}$ to $\alpha_{T}$. Step $T$ refers to the last noising step, when the sample is indistinguishable from pure noise, such that $\mathbf{m}_{T}\approx\mathcal{N}(\mathbf{0},\mathbf{I}_{N_{c}})$. Denoising refers to the opposite process, where noise is progressively subtracted from a perturbed geomodel until a clean (denoised) model is obtained. A neural network with learnable parameters $\theta$ (typically a U-net, (Ronneberger et al., 2015)) is trained to predict the noise $\bm{\epsilon}_{\theta}(\mathbf{m}_{t},t)$ given a randomly perturbed input $\mathbf{m}_{t}$ at each step $t$. As such, the training loss function is
@@
The specific DM method described above, and Eq. 3 in particular, correspond to the denoising diffusion implicit model, DDIM (Song et al., 2021a). This model differs from the original denoising diffusion probabilistic model, DDPM (Ho et al., 2020). As discussed in Di Federico and Durlofsky (2025), the DDIM variant is advantageous for our geomodeling application because it enables deterministic sampling in fewer steps than the DDPM. Only the DDIM will be considered throughout the rest of this paper.
Standard DMs operate in a multi-Gaussian latent space of the same dimension as the geomodel space, i.e., there is no dimension reduction. This makes the extension to large geomodels computationally expensive and provides no reduction in the number of variables to be determined during history matching. In LDMs, by contrast, noising and denoising are accomplished by a U-net in a reduced-dimension latent space. A variational autoencoder (VAE) performs the mapping between the latent space and the full geomodel space. We will now discuss the LDM procedure for 3D geomodel parameterization. Our modifications to the 2D-LDM method presented in Di Federico and Durlofsky (2025) include the use of 3D neural network layers, as well as a new perceptual loss term during VAE training.
@@
The 3D-LDM utilizes a latent variable $\bm{\xi}\in\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}$, where $n_{x}=N_{x}/f$, $n_{y}=N_{y}/f$, $n_{z}=N_{z}/f$, with $f$ being the downsampling ratio (Rombach et al., 2022). The overall dimension reduction ratio from the geomodel to latent space is $f^{3}$. The latent space, although lower-dimensional, is still 3D. The VAE comprises an encoder, such that $\mathcal{E}(\mathbf{m}_{0}):\mathbb{R}^{N_{x}\times N_{y}\times N_{z}}\rightarrow\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}$, and a decoder, $\mathcal{D}(\bm{\xi}_{0}):\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}\rightarrow\mathbb{R}^{N_{x}\times N_{y}\times N_{z}}$. Note the use of the subscript 0, referring to the encoder operating only on a clean geomodel, $\mathbf{m}_{0}$, and the decoder only on a fully denoised latent variable, $\bm{\xi}_{0}$. The VAE, with parameters $\theta$, is trained on the weighted loss function
@@
The perceptual loss addresses some of the limitations of pixel-wise losses (like the reconstruction loss), thus ensuring that decoded models (and, by extension, generated models) are perceptually similar to training models. Established pretrained networks used in this context include VGG16 (Simonyan and Zisserman, 2015), SqueezeNet (Iandola et al., 2016), AlexNet (Krizhevsky et al., 2012), and ResNet (He et al., 2016). For this study, we use ResNet.
A schematic of the VAE training procedure and a high-level overview of its architecture are presented in Figure 2a. The input model $\mathbf{m}_{0}$ is encoded into the latent variable $\bm{\xi}$, and decoded back into its reconstruction $\hat{\mathbf{m}}_{0}$. Reconstruction, perceptual, and hard data loss are computed as a function of $\mathbf{m}_{0}$ and $\hat{\mathbf{m}}_{0}$, while the K-L loss involves only $\bm{\xi}$. A 80-10-10% split for the training, validation, and testing sets is applied. Training requires approximately 12 h (2500 epochs) on one NVIDIA V100 GPU. We apply a learning rate of $10^{-4}$ with the Adam optimizer (Kingma and Ba, 2015), and a batch size of 16. Following Rombach et al. (2022), we set $\lambda_{\text{KL}}=10^{-6}$. Based on numerical experimentation, we specify $\lambda_{\text{h}}=10^{-2}$ and $\lambda_{\text{perc}}=10^{-3}$.
@@
Figure 2b depicts the U-net training procedure. The encoded geomodel, $\bm{\xi}_{0}$, which here represents a “clean” latent variable, is perturbed by a scheduler-added noise at a random discrete step $t$, using Eq. 7, to obtain $\bm{\xi}_{t}$. The U-net is then used to predict this noise from the inputs $\bm{\xi}_{t}$ and $t$, as $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$, and finally to compute $L_{\text{U-net}}$.

### v2

# 3D latent diffusion models for parameterizing and history matching facies systems under hierarchical uncertainty

@@ Abstract
Geological parameterization procedures entail the mapping of a high-dimensional geomodel to a low-dimensional latent variable. These parameterizations can be very useful for history matching because the number of variables to be calibrated is greatly reduced, and the mapping can be constructed such that geological realism is automatically preserved. In this work, a parameterization method based on generative latent diffusion models (LDMs) is developed for 3D channel-levee-mud systems. Geomodels with variable geological metaparameters (also referred to as hyperparameters or scenario parameters), specifically mud fraction, channel orientation, and channel width, are considered. A perceptual loss term is included during training to improve geological realism. For any set of scenario parameters, an (essentially) infinite number of realizations can be generated, so our LDM parameterizes over a very wide model space. New realizations constructed using the LDM procedure are shown to closely resemble reference geomodels, both visually and in terms of one- and two-point spatial statistics. Flow response distributions, for a specified set of injection and production wells, are also shown to be in close agreement between the two sets of models. The parameterization method is applied for ensemble-based history matching, with model updates performed in the LDM latent space, for cases involving uncertain scenario parameters. For three synthetic true models, we observe clear uncertainty reduction in both production forecasts and geological scenario parameters. The overall method is additionally shown to provide posterior geomodels consistent with the synthetic true model in each case.
@@ 1 Introduction
In this work, we introduce a 3D deep learning-based latent diffusion model for the parameterization of facies-based geological systems. We consider relatively large channel-levee-mud geomodels, and introduce a parameterization that can treat variable continuous geological metaparameters/scenario parameters. This represents a significant extension of our previous latent diffusion model where we considered 2D, fixed-metaparameter systems (DIFEDERICO2025105755). We demonstrate the usefulness of the parameterization method by performing history matching with an ensemble smoother with multiple data assimilation (ESMDA) procedure (EMERICK20133) operating on latent space variables.
@@
In recent years, a number of deep learning (DL) geomodel parameterizations have been introduced. These have been shown to perform better than more mathematically tractable methods, such as those based on principal component analysis (e.g., Sarma2008) or discrete cosine transform (Jafarpour), which are not always able to reproduce complex geological patterns (Chan_2019). DL methods learn the mapping between geological model space and latent space from a training dataset, and are then able to generate new realizations that exhibit the same spatial features. Architectures for this application include CNN-PCA (LIU2021104676), variational autoencoders (VAEs) (Canchumuni_2019; mo_caae; grana_co2), and generative adversarial networks (GANs) (Laloy_2018; chan2019parametrization; Canchumuni_2020; HU2023105290; Feng2024; Merzoug2025). Existing DL parameterizations have proved successful both for standalone geomodeling and history matching for various subsurface flow applications. These include, for example, oil and gas production (Tang_2021), groundwater flow (Song_Mukerji_Hou_2021), and geological carbon storage (BAO2024213294). These methods do, however, have some limitations. Namely, VAEs may provide overly smoothed or discontinuous geological patterns (Canchumuni_2019; chan2019parametrization; abbate), while GANs often require computationally intensive (potentially unstable) adversarial training in conjunction with careful hyperparameter tuning. Improved treatments have been presented, including VAE-GAN hybrids (ZHAO2023105239; FENG2024104833), progressive training (Song_Mukerji_Hou_2021), StyleGAN (stylegan), and the use of the Wasserstein loss (Chan_2019). Although these developments are promising, the limitations of existing methods have motivated the application of the more recent diffusion models (DMs), which currently represent the state-of-the-art in generative modeling.
@@
DMs, first introduced by ho2020denoising, are a class of generative models trained to learn a denoising process. During denoising, a neural network is used to predict the noise contained in a sample, which is then iteratively subtracted for a number of steps. In this way, the DM transforms an initial noise input into a clean, “denoised” image. To accomplish this, the DM learns the mapping between a Gaussian noise realization (latent space) and the corresponding geological model (model space). DMs have been applied for geomodeling purposes in a few studies, and they have been shown to provide high-quality unconditional and conditional realizations for complex geological patterns. These include fluvial channel (mosser; Xu2024DiffSimDD; DIFEDERICO2025105755), Gaussian (alnasser), and discrete fracture systems (fractures_ldm).
@@
Among DMs, latent diffusion models (LDMs, introduced by rombach2022highresolution), appear to be a particularly promising method for geomodel generation and history matching in subsurface flow applications. These methods achieve computational efficiency and scalability by performing the diffusion process in a lower-dimensional latent space rather than in the pixel or voxel space. lee2023latent and Ovanger2025 used LDMs to generate realistic conditional realizations of 2D multifacies marine depositional environments. These studies highlighted the superior performance of diffusion-based generation relative to GANs. In fractures_ldm, an LDM was used to parameterize 2D discrete fracture networks (DFNs). LDM generated models exhibited visual improvement over those generated using VAEs, and they preserved property distributions in agreement with reference models. LDMs have also been used for direct generation or post-processing of simulation forecasts. For example, zhan applied an LDM for 2D binary facies models with Gaussian permeability distributions within the two facies. In that study an LDM was trained to generate facies structures conditioned on observed pressure and concentration fields, and another LDM with the same architecture was trained to perform the opposite task. exxon developed an LDM model to post-process upscaled simulation results to provide fine-scale pressure and saturation maps for a 2D waterflooding problem.
@@
To date, many subsurface modeling applications involving DL parameterizations – and in particular most DM/LDM implementations – have been for 2D or relatively small 3D domains, containing fewer than $\sim$100,000 cells. Given their powerful generative capabilities and efficient representation of high-dimensional spaces, the extension and application of these methods to larger 3D models is of great interest. In addition, with the exception of the work in Song2021_multiscenario for 2D geomodeling with GANs, the global features characterizing the geological scenario are typically fixed in studies involving flow and history matching. This could represent a key limitation because the scenario parameters themselves are uncertain in practical applications, and this should be considered during history matching.
@@
In this work, we address both of these important issues by (1) developing a 3D-LDM applicable over a range of continuous scenario parameters and (2) applying this treatment to history matching for cases where these parameters are uncertain. The 3D models considered in this work, of grid dimensions 128 $\times$ 128 $\times$ 32 (total of 524,288 cells), are larger than many of the models treated in previous studies. In addition, the three-facies geomodels entail uncertain scenario parameters, namely mud fraction, channel width, and channel orientation.
@@
New samples generated by the 3D-LDM method are shown to exhibit high visual quality in terms of geological features, along with spatial statistics that are consistent with (reference) training samples from Petrel geomodeling software (Schlumberger). Significant computational savings are achieved in the geomodeling process. Flow statistics for prior models confirm that the parameterization provides median results, and the correct amount of variability, relative to reference results over a range of metaparameters. History matching is then performed for waterflooding problems using the 3D-LDM representation in combination with an ESMDA procedure (EMERICK20133). Results for cases involving synthetic true models corresponding to different scenario parameters are presented.
@@
This paper proceeds as follows. In Section 2, we describe the channel-levee-mud geomodels considered in this work. In Section 3, after reviewing the basics of DMs/LDMs and their use for geomodeling, we describe the 3D-LDM parameterization method and the new treatments introduced in this work. Next, in Section 4, we display LDM geomodel realizations and compare spatial and flow statistics for an ensemble of LDM generated models to those for reference Petrel models. Latent-space history matching results, for different synthetic true models, are provided in Section 5. Finally, in Section 6, we summarize this work and suggest directions for future research. Additional algorithmic details are provided in the appendices.
2 3D fluvial channel systems with variable scenario parameters
@@
In this work, we define a geological scenario as a realization of a set of continuous geological scenario parameters/metaparameters that characterize the depositional environment. This approach allows for variation within a general geological concept. For example, in the systems considered here, the metaparameters quantify the geometries of channel bodies in fluvial channel systems. It is important to emphasize that, with the scenario parameters considered here, we are not attempting to treat widely different depositional environments corresponding to distinct geological concepts. For example, these parameters are not meant to characterize settings ranging from fluvial channel systems to eolian sand-dune systems to fractured carbonate systems. Nonetheless, even within a particular geological concept, an infinite number of realizations can be constructed for any set of scenario parameters. In this sense the geomodel uncertainty is hierarchical, as we have uncertain higher-level scenario parameters and, for each set of parameters, uncertain geobody realizations. We will refer to this as hierarchical uncertainty rather than multiscenario uncertainty, since the latter term could also refer to cases with widely different depositional environments (which are not treated here).
@@
We can express a particular geomodel $\mathbf{m}$ as $\mathbf{m}(\mathbf{s},\mathbf{r})$, where $\mathbf{s}$ denotes the set of scenario parameters and $\mathbf{r}$ indicates the particular realization. In our case, $\mathbf{r}$ corresponds to the cell-by-cell facies indicator. The fluvial channel systems considered here contain channel (high permeability), levee (intermediate permeability), and mud/shale (low permeability) facies. The levee facies surrounds the extensive channel bodies, with both deposited within a mud background. Mud, levee, and channel facies are identified in the geomodels by cell values of 0, 0.5, and 1, respectively. The geological scenario parameters $\mathbf{s}$ include mud fraction (denoted $f_{m}$), channel orientation ($\theta_{\rm ch}$), and channel width ($w_{\rm ch}$). Thus, $\mathbf{s}=[f_{m},\theta_{\rm ch},w_{\rm ch}]$. The geomodel realizations used for training, and as the reference against which the LDM models will be compared, are constructed using the Petrel geomodeling software (Schlumberger). Object-based modeling (OBM) is applied.
@@
Mud fraction, $f_{m}$ 0.72 0.87
Channel orientation (avg. angle), $\theta_{\rm ch}$ 30 60
Channel width (avg. # of blocks), $w_{\rm ch}$ 4 6
@@
Figure 1: (a) Conditioning locations for all geomodels and well labels for (later) flow simulations. (b–f) Representative realizations corresponding to different scenario parameters from the Petrel-generated training set (scenario parameters are given for each geomodel, channel width in units of grid cells).
@@ 3 Latent diffusion models for representing geological systems
In this section, we first present an overview of the general DM (ho2020denoising) and LDM (rombach2022highresolution) procedures. Our 3D-LDM method for geological parameterization with variable metaparameters is then described. From a geomodeling perspective, scenario parameters act as a form of conditioning to global features, as opposed to conditioning to particular values at specific spatial locations, i.e., hard data. Strategies to implement scenario conditioning in a deep learning parameterization context include the use of additional input channels or separate parameter-encoding pipelines. These approaches enable direct control on the generation process, and have proved effective for geomodeling purposes (Song_Mukerji_Hou_2021). In this study, however, we adopt a simpler approach, where the method is trained to parameterize for a range of scenario parameters without additional constraints or input features.
@@ 3.1 Diffusion model (DM)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\mathbf{I}_{N_{c}}$ is the identity matrix. The scheduler defines the values of the perturbation through the $\alpha_{s}$ parameter at each step. This is typically a linearly decreasing function from $\alpha_{1}$ to $\alpha_{T}$. Step $T$ refers to the last noising step, when the sample is indistinguishable from pure noise, such that $\mathbf{m}_{T}\approx\mathcal{N}(\mathbf{0},\mathbf{I}_{N_{c}})$. Denoising refers to the opposite process, where noise is progressively subtracted from a perturbed geomodel until a clean (denoised) model is obtained. A neural network with learnable parameters $\theta$ (typically a U-net, (unet)) is trained to predict the noise $\bm{\epsilon}_{\theta}(\mathbf{m}_{t},t)$ given a randomly perturbed input $\mathbf{m}_{t}$ at each step $t$. As such, the training loss function is
@@
The specific DM method described above, and Eq. 3 in particular, correspond to the denoising diffusion implicit model, DDIM (song2021denoising). This model differs from the original denoising diffusion probabilistic model, DDPM (ho2020denoising). As discussed in DIFEDERICO2025105755, the DDIM variant is advantageous for our geomodeling application because it enables deterministic sampling in fewer steps than the DDPM. Only the DDIM will be considered throughout the rest of this paper.
Standard DMs operate in a multi-Gaussian latent space of the same dimension as the geomodel space, i.e., there is no dimension reduction. This makes the extension to large geomodels computationally expensive and provides no reduction in the number of variables to be determined during history matching. In LDMs, by contrast, noising and denoising are accomplished by a U-net in a reduced-dimension latent space. A variational autoencoder (VAE) performs the mapping between the latent space and the full geomodel space. We will now discuss the LDM procedure for 3D geomodel parameterization. Our modifications to the 2D-LDM method presented in DIFEDERICO2025105755 include the use of 3D neural network layers, as well as a new perceptual loss term during VAE training.
@@
The 3D-LDM utilizes a latent variable $\bm{\xi}\in\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}$, where $n_{x}=N_{x}/f$, $n_{y}=N_{y}/f$, $n_{z}=N_{z}/f$, with $f$ being the downsampling ratio (rombach2022highresolution). The overall dimension reduction ratio from the geomodel to latent space is $f^{3}$. The latent space, although lower-dimensional, is still 3D. The VAE comprises an encoder, such that $\mathcal{E}(\mathbf{m}_{0}):\mathbb{R}^{N_{x}\times N_{y}\times N_{z}}\rightarrow\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}$, and a decoder, $\mathcal{D}(\bm{\xi}_{0}):\mathbb{R}^{n_{x}\times n_{y}\times n_{z}}\rightarrow\mathbb{R}^{N_{x}\times N_{y}\times N_{z}}$. Note the use of the subscript 0, referring to the encoder operating only on a clean geomodel, $\mathbf{m}_{0}$, and the decoder only on a fully denoised latent variable, $\bm{\xi}_{0}$. The encoder outputs the parameters of a Gaussian distribution, $(\mathcal{E}_{\mu},\mathcal{E}_{\sigma})$, and the latent variable is sampled via the reparameterization $\bm{\xi}_{0}=\mathcal{E}_{\mu}+\mathcal{E}_{\sigma}\cdot\bm{\epsilon},\quad\text{where}\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n_{c}})$. Overall, we denote $\bm{\xi}_{0}=\mathcal{E}(\mathbf{m}_{0})$.
The VAE, with parameters $\theta$, is trained on the weighted loss function
@@
The perceptual loss addresses some of the limitations of pixel-wise losses (like the reconstruction loss), thus ensuring that decoded models (and, by extension, generated models) are perceptually similar to training models. Established pretrained networks used in this context include VGG16 (vgg), SqueezeNet (squeezenet2016), AlexNet (alexnet), and ResNet (resnet50). For this study, we use ResNet.
A schematic of the VAE training procedure and a high-level overview of its architecture are presented in Figure 2a. The input model $\mathbf{m}_{0}$ is encoded into the latent variable $\bm{\xi}$, and decoded back into its reconstruction $\hat{\mathbf{m}}_{0}$. Reconstruction, perceptual, and hard data loss are computed as a function of $\mathbf{m}_{0}$ and $\hat{\mathbf{m}}_{0}$, while the K-L loss involves only $\bm{\xi}$. Pseudocode is provided in Algorithm 1 in Appendix 3D-LDM algorithms. A 80-10-10% split for the training, validation, and testing sets is applied. Training requires approximately 12 h (2500 epochs) on one NVIDIA A100 GPU. We apply a learning rate of $10^{-4}$ with the Adam optimizer (adam), and a batch size of 16. Following rombach2022highresolution, we set $\lambda_{\text{KL}}=10^{-6}$. Based on numerical experimentation, we specify $\lambda_{\text{h}}=10^{-2}$ and $\lambda_{\text{perc}}=10^{-3}$.
@@
Figure 2b depicts the U-net training procedure. The encoded geomodel, $\bm{\xi}_{0}$, which here represents a “clean” latent variable, is perturbed by a scheduler-added noise at a random discrete step $t$, using Eq. 7, to obtain $\bm{\xi}_{t}$. The U-net is then used to predict this noise from the inputs $\bm{\xi}_{t}$ and $t$, as $\bm{\epsilon}_{\theta}(\bm{\xi}_{t},t)$, and finally to compute $L_{\text{U-net}}$. Pseudocode for the U-net training is given in Algorithm 2 in Appendix 3D-LDM algorithms.